# Normal generators for mapping class groups

### v1

@@ 5. Asymptotic translation lengths on curve graphs
In the rest of the article, let $S=S_{g}$ be a closed surface of genus $g\geq 2$. There is another metric space on which the mapping class group $\operatorname{Mod}(S)$ acts by isometries, the curve graph. In this section, we discuss the asymptotic translation length on the curve graph and its relation to normal generation.
@@
The curve graph $\mathcal{C}(S)$ of $S$ was first introduced by Harvey [Har81]. It is a graph whose vertices are isotopy classes of essential simple closed curves on $S$, and two vertices are connected by an edge if they have disjoint representatives. We equipped with a simplicial metric $d_{\mathcal{C}}$ on the curve graph. Some geometric properties were studied by Masur and Minsky [MM99].
Theorem 5.1 (Masur-Minsky).
@@
It is clear from the definition that $\operatorname{Mod}(S)$ isometrically on $\mathcal{C}(S)$. We use the notation
@@
for the asymptotic translation length on $\mathcal{C}(S)$. It was also shown in [MM99, Proposition 3.6] that if $f\in\operatorname{Mod}(S)$ is pseudo-Anosov, then
@@
It is easy to see that periodic mapping classes and reducible mapping classes have zero asymptotic translation lengths on the curve graph.
Hence it is natural to ask whether pseudo-Anosovs with small $\ell_{\mathcal{C}}$ are normal generators, analogous to Theorem 4.2. This question is widely open, and this section is devoted to introduce some work towards this (cf. [BKSW23, Question 1.2]).
@@
[…]
@@
On the other hand, things are very different when we consider asymptotic translation lengths on curve graphs. Analogous to (4.1) and (5.1), we define
[…]
@@
In the viewpoint of (5.3), we formulate the question pertaining to the relation between translation lengths on curve graphs and noraml generation as follows:
@@
One can also ask whether Torelli group is the only obstruction for pseudo-Anosov with small $\ell_{\mathcal{C}}$ to be a normal generator:
@@
Both questions above are still open. However, in our joint work with Wu [BKW21a], we gave an upper bound for $C$ in Question 5.3, if its answer is yes.
Theorem 5.4 (Baik-Kim-Wu).
@@
The rest of this article is devoted to explain how we constructed a sequence in the above theorem. An idea to construct pseudo-Anosov mapping classes that do not normally generate mapping class groups is considering finite covers and taking the lifts of a fixed pseudo-Anosov mapping class. Such lifts possess periodic behavior, from which we deduce that they are not normal generators. Moreover, as the genus gets bigger, more simple closed curves have common disjoint simple closed curves, which would make the distance in curve graphs smaller. It would be a good exercise to modify the construction to obtain the upper bound of $L_{\mathcal{C}}(\mathcal{I}_{g})$ in (5.3). See the work of Baik and Shin [BS20] for the proof of the precise asymptote of $L_{\mathcal{C}}(\mathcal{I}_{g})$.
@@
Let $\alpha$ be a non-separating simple closed curve on the closed surface $S_{2}$ of genus $2$. Cutting $S_{2}$ by $\alpha$, and gluing $g$ copies of the resulting surface along copies of $\alpha$ in a cyclic way, we obtain the closed surface $S_{g+1}$ of genus $g+1$. This gives the finite cyclic cover $p_{g+1}$ of degree $g$, as described in Figure 4.
$\alpha$$p_{g+1}$
@@
The covering $p_{g+1}$ can also be defined algebraically. Recall the algebraic intersection number $\hat{i}(\cdot,\cdot)$. Considering the composition
@@
$[f^{-1}(\alpha)]=[T_{\varphi\alpha}T_{\varphi\beta}T_{\beta}^{-1}\alpha]=[T_{\varphi\alpha}\alpha]=[\varphi T_{\alpha}\varphi^{-1}\alpha]=[\alpha]\in H_{1}(S_{2})$
@@
Let us show that $\tilde{f}$ is not a Torelli element of $\operatorname{Mod}(S_{g+1})$. Let $\eta$ and $\tilde{\eta}$ be simple closed curves as in Figure 7. We then have $[p(\tilde{\eta})]=g[\eta]\in H_{1}(S_{2})$. Since $f=T_{\beta}T_{\varphi\beta}^{-1}T_{\varphi\alpha}^{-1}$ and $T_{\beta}T_{\varphi\beta}^{-1}$ is Torelli, this implies
$[f(\eta)]=[T_{\varphi\alpha}^{-1}\alpha]=[\varphi T_{\alpha}^{-1}\varphi^{-1}\eta].$
@@
$[f(\eta)]=[T_{\alpha}^{-1}\eta]\neq[\eta].$
This yields $[\tilde{f}(\tilde{\eta})]\neq[\tilde{\eta}]$, and therefore $\tilde{f}$ is not Torelli.
$\alpha$$\eta$$p$$\tilde{\eta}$
@@
As one can see from Figure 5, each component of $p^{-1}(\beta)$ is separating. Hence, $T_{p^{-1}(\beta)}$ is Torelli. In particular, $T_{p^{-1}(\beta)}$ acts trivially on $H_{1}(S_{g+1};\mathbb{Z}/g\mathbb{Z})$. In addition, any two components of $p^{-1}(\alpha)$ bound a subsurface, and hence they are homologous. Fixing a component $\tilde{\alpha}$ of $p^{-1}(\alpha)$, this implies that the action of $T_{p^{-1}(\alpha)}$ on $H_{1}(S_{g+1})$ is identical to the action of $T_{\tilde{\alpha}}^{g}$, which acts trivially on $H_{1}(S_{g+1};\mathbb{Z}/g\mathbb{Z})$. This finishes the proof of the claim.
@@
We label components of $S_{g+1}-p^{-1}(\alpha)$ by $X_{1},\cdots,X_{g}\subset S_{g+1}$ so that $X_{i}$ and $X_{i+1}$ are glued along one of their boundary components for all $1\leq i\leq g$, writing the index $i$ modulo $g$. We keep this convention throughout the section; in particular, $X_{0}=X_{g}$. We make the choice of $\tilde{\alpha}$ more explicit by setting $\tilde{\alpha}:=\partial X_{0}\cap\partial X_{1}$.
@@
$\alpha$$p$$\tilde{\alpha}$
Figure 8. Geometric intersection number of $\alpha$ and a curve on $S_{2}$ determines by which subsurfaces the multitwist of $\tilde{\alpha}$ along the preimage of the curve is trapped as in (5.4).
@@
Inductively, we have that for any $m\in\mathbb{N}$,
@@
Therefore, the estimate for $\ell_{\mathcal{C}}(\tilde{f})$ follows once we obtain the largest possible $m$. Recall that $\lambda=T_{\xi}\beta$ and $\varphi=T_{\lambda}T_{\beta}^{-1}$. From Figure 6, we have
@@
$i(\varphi\alpha,\alpha)$ $=i(T_{\lambda}\alpha,\alpha)=i(\lambda,\alpha)^{2}=144;$
$i(\varphi\beta,\alpha)$ $=i(T_{\lambda}\beta,\alpha)=i(\lambda,\beta)i(\lambda,\alpha)=432.$
@@ 6. Further questions
We finish the article by recording some further questions that have not yet been answered. Let $S_{g}$ be a closed surface of genus $g\geq 2$. We first recall the questions mentioned above:
@@
In this regard, one can also ask a weaker version of Question 6.1 focusing on pseudo-Anosov mapping classes whose asymptotic translation lengths on curve graphs are minimal in the mapping class group.
@@
The similar question for asymptotic translation lengths on Teichmüller spaces has the affirmative answer by Theorem 4.2 (Lanier-Margalit [LM22]) and Theorem 4.4 (Penner [Pen91]). By Theorem 4.4, we have
@@
Since Theorem 4.2 asserts that a pseudo-Anosov $f\in\operatorname{Mod}(S_{g})$ is a normal generator if $\ell_{\mathcal{T}}(f)\leq\frac{1}{2}\log 2$ for $g\geq 3$, this answers the $\ell_{\mathcal{T}}$-version of Question 6.3 affirmative.
We also remark that the difference between two asymptotes (6.1) and (6.2) explains the reason for having the genus in the upper bounds for $\ell_{\mathcal{C}}$ in Question 6.1 and Question 6.2.
Lanier-Margalit’s criterion for general mapping classes
The original Lanier-Margalit’s criterion (Theorem 4.2) is about pseudo-Anosov mapping classes. As stated in Theorem 4.6, it was extended in our joint work with Wu to partly pseudo-Anosov mapping classes, which include certain reducible elements in $\operatorname{Mod}(S_{g})$. In this regard, we ask for the largest subclass of $\operatorname{Mod}(S_{g})$ to which Lanier-Margalit’s criterion applies. We first ask whether the criterion applies to all reducible mapping classes.
@@
Is there an alternative characterization for the maximal subset of $\operatorname{Mod}(S_{g})$ to which Lanier-Margalit’s criterion applies?
@@
Let $V_{g}$ be a handlebody of genus $g\geq 2$. That is, the 3-manifold with boundary obtained by attaching the $g$ number of the 1-handle to the 3-ball. We identify $S_{g}$ with $\partial V_{g}$ and consider the following subgroup of $\operatorname{Mod}(S_{g})$, called the handlebody group:
@@
[…]
@@
Therefore, we apply Lanier-Margalit’s criterion (Theorem 4.2) and conclude that there are normal generators of $\operatorname{Mod}(S_{g})$ in $\mathcal{H}_{g}$ for all large enough $g$.
@@
Again, we refer to such an element $f\in\mathcal{H}_{g}$ as a normal generator of $\mathcal{H}_{g}$. It might be natural to expect that there are normal generators of $\mathcal{H}_{g}$ among normal generators of $\operatorname{Mod}(S_{g})$ in $\mathcal{H}_{g}$, which exist as observed above. From the viewpoint of Lanier-Margalit’s criterion, we also ask whether the small asymptotic translation length $\ell_{\mathcal{T}}$ implies a normal generation of $\mathcal{H}_{g}$.
@@
    $[\operatorname{Mod}(S_{g}),\operatorname{Mod}(S_{g})]\leq\langle\langle f\rangle\rangle.$
@@
    $[\operatorname{Mod}(S_{g}),\operatorname{Mod}(S_{g})]=\operatorname{Mod}(S_{g}).$
@@
Does the well-suited criterion hold for $\mathcal{H}_{g}$? More precisely, is there a graph $N_{f}$ associated to each $f\in\mathcal{H}_{g}$ so that the connectedness of $N_{f}$ implies that $[\mathcal{H}_{g},\mathcal{H}_{g}]$ is contained in the normal closure of $f$ in $\mathcal{H}_{g}$?

### v2

@@ 5. Asymptotic translation lengths on curve graphs
In the rest of the chapter, let $S=S_{g}$ be a closed surface of genus $g\geq 2$. There is another metric space on which the mapping class group $\operatorname{Mod}(S)$ acts by isometries, namely, the curve graph. In this section, we discuss the asymptotic translation length on the curve graph and its relation to normal generation.
@@
The curve graph $\mathcal{C}(S)$ of $S$ was first introduced by Harvey [HAR81]. It is a graph whose vertices are isotopy classes of essential simple closed curves on $S$, and where two vertices are connected by an edge if they have disjoint representatives. We equip the curve graph with a simplicial metric $d_{\mathcal{C}}$. Some geometric properties of $\mathcal{C}(S)$ were studied by Masur and Minsky [MM99].
Theorem 5.1 (Masur–Minsky).
@@
Here, a geodesic metric space is called Gromov hyperbolic if there exists $\delta>0$ so that every geodesic triangle is $\delta$-thin, i.e., the $\delta$-neighborhood of one side contains the other two sides.
It is clear from the definition that $\operatorname{Mod}(S)$ acts isometrically on $\mathcal{C}(S)$. We use the notation
@@
for the asymptotic translation length on $\mathcal{C}(S)$. Since there exists a coarsely $\operatorname{Mod}(S)$-equivariant and coarsely Lipschitz map $\mathcal{T}(S)\to\mathcal{C}(S)$ [MM99], there exists $c=c(S)>0$ such that $\ell_{\mathcal{C}}(f)\leq c\cdot\ell_{\mathcal{T}}(f)$ for all $f\in\operatorname{Mod}(S)$. A more effective version of this estimate is obtained in [GHK+13]. On the other hand, $\ell_{\mathcal{C}}$ and $\ell_{\mathcal{T}}$ are not close to each other in general [BAD25].
The Gromov hyperbolicity of $\mathcal{C}(S)$ provides more tools to study the dynamics of $\operatorname{Mod}(S)$ on $\mathcal{C}(S)$. It was also shown in [MM99, Proposition 3.6] that if $f\in\operatorname{Mod}(S)$ is pseudo-Anosov, then
@@
It is easy to see that periodic mapping classes and reducible mapping classes have zero asymptotic translation lengths on the curve graph since they have powers fixing a vertex in $\mathcal{C}(S)$. In contrast, note that reducible mapping classes can have positive asymptotic translation lengths on $\mathcal{T}(S)$.
Hence it is natural to ask whether pseudo-Anosovs with small $\ell_{\mathcal{C}}$ are normal generators, in analogy with to Theorem 4.2. This question is wide open, and this section is devoted to introduce some work towards it (cf. [BKS+23, Question 1.2]).
@@
(5.1) $L_{\mathcal{T}}(\mathcal{I}_{g}):=\inf\{\ell_{\mathcal{T}}(f):f\in\mathcal{I}_{g}\text{ is pseudo-Anosov}\},$
@@
On the other hand, things are very different when we consider asymptotic translation lengths on curve graphs. In analogy with (4.1) and (5.1), we define
$L_{\mathcal{C}}(g)$ $=\inf\{\ell_{\mathcal{C}}(f):f\in\operatorname{Mod}(S_{g})\text{ is pseudo-Anosov}\};$
$L_{\mathcal{C}}(\mathcal{I}_{g})$ $=\inf\{\ell_{\mathcal{C}}(f):f\in\mathcal{I}_{g}\text{ is pseudo-Anosov}\}.$
@@
In view of (5.3), we formulate the following question pertaining to the relation between translation lengths on curve graphs and normal generation as follows:
@@
Does there exist a constant $C>0$ such that if $f\in\operatorname{Mod}(S_{g})$ is pseudo-Anosov with $\ell_{\mathcal{C}}(f)\leq\frac{C}{g}$, then $f$ is a normal generator?
One can also ask whether the Torelli group is the only obstruction for a pseudo-Anosov with small $\ell_{\mathcal{C}}$ to be a normal generator:
@@
Both questions are still open. However, in our joint work with Wu [BKW24], we gave an upper bound for $C$ in Question 5.3, if its answer is yes.
Theorem 5.4 (Baik–Kim–Wu).
@@
The quantities 577 and 578 arise from intersection numbers between curves whose associated Dehn twists are used in the construction of $f_{g}$. The rest of this chapter is devoted to explain how we constructed a sequence in the above theorem. An idea to construct pseudo-Anosov mapping classes that do not normally generate mapping class groups is considering finite covers and taking the lifts of a fixed pseudo-Anosov mapping class. Such lifts possess a periodic behavior, from which we deduce that they are not normal generators. Moreover, as the genus gets bigger, more simple closed curves have common disjoint simple closed curves, which would make the distance in the curve graphs smaller. It would be a good exercise to modify the construction to obtain the upper bound of $L_{\mathcal{C}}(\mathcal{I}_{g})$ in (5.3). See the work of Baik and Shin [BS20] for the proof of the precise asymptote of $L_{\mathcal{C}}(\mathcal{I}_{g})$.
@@
Let $\alpha$ be a non-separating simple closed curve on the closed surface $S_{2}$ of genus $2$. Cutting $S_{2}$ along $\alpha$ and gluing $g$ copies of the resulting surface along copies of $\alpha$ in a cyclic way, we obtain the closed surface $S_{g+1}$ of genus $g+1$. This gives the finite cyclic cover $p_{g+1}$ of degree $g$, as described in Figure 4.
[…]
@@
[…]
@@
$f^{-1}(\alpha)]=[T_{\varphi\alpha}T_{\varphi\beta}T_{\beta}^{-1}\alpha]=[T_{\varphi\alpha}\alpha]=[\varphi T_{\alpha}\varphi^{-1}\alpha]=[\alpha]\in H_{1}(S_{2}).$
@@
Let us show that $\tilde{f}$ is not a Torelli element of $\operatorname{Mod}(S_{g+1})$. Let $\eta$ and $\tilde{\eta}$ be simple closed curves as in Figure 7. We then have $p(\tilde{\eta})]=g[\eta]\in H_{1}(S_{2})$. Since $f=T_{\beta}T_{\varphi\beta}^{-1}T_{\varphi\alpha}^{-1}$ and $T_{\beta}T_{\varphi\beta}^{-1}$ is Torelli, this implies
$f(\eta)]=[T_{\varphi\alpha}^{-1}\alpha]=[\varphi T_{\alpha}^{-1}\varphi^{-1}\eta].$
@@
$f(\eta)]=[T_{\alpha}^{-1}\eta]\neq[\eta].$
This yields $\tilde{f}(\tilde{\eta})]\neq[\tilde{\eta}$], and therefore $\tilde{f}$ is not Torelli.
[…]
@@
As one can see from Figure 5, each component of $p^{-1}(\beta)$ is separating. Hence, $T_{p^{-1}(\beta)}$ is Torelli. In particular, $T_{p^{-1}(\beta)}$ acts trivially on $H_{1}(S_{g+1};\mathbb{Z}/g\mathbb{Z})$. In addition, any two components of $p^{-1}(\alpha)$ bound a subsurface, hence they are homologous. Fixing a component $\tilde{\alpha}$ of $p^{-1}(\alpha)$, this implies that the action of $T_{p^{-1}(\alpha)}$ on $H_{1}(S_{g+1})$ is identical to the action of $T_{\tilde{\alpha}}^{g}$, which acts trivially on $H_{1}(S_{g+1};\mathbb{Z}/g\mathbb{Z})$. This finishes the proof of the claim.
@@
We label the components of $S_{g+1}-p^{-1}(\alpha)$ by $X_{1},\cdots,X_{g}\subset S_{g+1}$ so that $X_{i}$ and $X_{i+1}$ are glued along one of their boundary components for all $1\leq i\leq g$, writing the index $i$ modulo $g$. We keep this convention throughout the section; in particular, $X_{0}=X_{g}$. We make the choice of $\tilde{\alpha}$ more explicit by setting $\tilde{\alpha}:=\partial X_{0}\cap\partial X_{1}$.
@@
[…]
Figure 8. The geometric intersection number of $\alpha$ and a curve on $S_{2}$ determines by which subsurfaces the multitwist of $\tilde{\alpha}$ along the preimage of the curve is trapped as in (5.4).
@@
Inductively, we have, for any $m\in\mathbb{N}$,
@@
[…]
@@
[…]
@@ 6. Further questions
We end this chapter by recording some further questions that have not yet been answered. Let $S_{g}$ be a closed surface of genus $g\geq 2$. We first recall the questions mentioned above:
@@
Hence, as a first step towards Question 6.1, one can also ask the following weaker version focusing on pseudo-Anosov mapping classes whose asymptotic translation lengths on curve graphs are minimal in the mapping class group.
@@
A similar question for asymptotic translation lengths on Teichmüller spaces has an affirmative answer by Theorem 4.2 (Lanier–Margalit [LM22]) and Theorem 4.4 (Penner [PEN91]). By Theorem 4.4, we have
@@
Since Theorem 4.2 asserts that a pseudo-Anosov $f\in\operatorname{Mod}(S_{g})$ is a normal generator if $\ell_{\mathcal{T}}(f)\leq\frac{1}{2}\log 2$ for $g\geq 3$, this answers the $\ell_{\mathcal{T}}$-version of Question 6.3 affirmatively.
We also remark that the difference between the two asymptotes (6.1) and (6.2) explains the reason for having the genus in the upper bounds for $\ell_{\mathcal{C}}$ in Question 6.1 and Question 6.2.
Lanier–Margalit’s criterion for general mapping classes
The original Lanier–Margalit criterion (Theorem 4.2) is about pseudo-Anosov mapping classes. As stated in Theorem 4.6, it was extended in our joint work with Wu to partly pseudo-Anosov mapping classes, which include certain reducible elements in $\operatorname{Mod}(S_{g})$. While this extension still requires mapping classes to have pseudo-Anosov dynamics on some subsurfaces, it is also true that Dehn twists along non-separating essential simple closed curves are also normal generators ([LIC64], [MUM67]). Such mapping classes do not exhibit any pseudo-Anosov dynamics on any subsurface, and have zero translation lengths on $\mathcal{T}(S_{g})$. In this regard, we ask for the largest subclass of $\operatorname{Mod}(S_{g})$ to which Lanier–Margalit’s criterion applies. We first ask whether the criterion applies to all reducible mapping classes.
@@
Is there an alternative characterization for the maximal subset of $\operatorname{Mod}(S_{g})$ to which Lanier–Margalit’s criterion applies?
@@
While the whole discussion so far has been about the normal generation of the full mapping class group $\operatorname{Mod}(S_{g})$. We now change gears and consider a slightly more general question: whether a given subgroup can also be normally generated, or whether there exists a neat description of its normal generators. We mainly discuss a concrete subgroup of $\operatorname{Mod}(S_{g})$, the handlebody group, which is an interesting object in itself and has been studied from various points of view. One can also refer to ([DGO17], [CMM21]) for related discussions on free subgroups.
Let $V_{g}$ be a handlebody of genus $g\geq 2$, that is, the 3-manifold with boundary obtained by attaching $g$ 1-handles to the 3-ball. We identify $S_{g}$ with $\partial V_{g}$ and consider the following subgroup of $\operatorname{Mod}(S_{g})$, called the handlebody group:
@@
(6.3) $L_{\mathcal{T}}(\mathcal{H}_{g}):=\inf\{\ell_{\mathcal{T}}(f):f\in\mathcal{H}_{g}\text{ is pseudo-Anosov}\}.$
@@
Therefore, we apply Lanier–Margalit’s criterion (Theorem 4.2) and conclude that there are normal generators of $\operatorname{Mod}(S_{g})$ in $\mathcal{H}_{g}$ for all large enough $g$.
@@
Again, we refer to such an element $f\in\mathcal{H}_{g}$ as a normal generator of $\mathcal{H}_{g}$. It might be natural to expect that there are normal generators of $\mathcal{H}_{g}$ among normal generators of $\operatorname{Mod}(S_{g})$ in $\mathcal{H}_{g}$, which exist as observed above. From the viewpoint of Lanier–Margalit’s criterion, we also ask whether the small asymptotic translation length $\ell_{\mathcal{T}}$ implies a normal generation of $\mathcal{H}_{g}$.
@@
    $\operatorname{Mod}(S_{g}),\operatorname{Mod}(S_{g})]\leq\langle\langle f\rangle\rangle.$
@@
    $\operatorname{Mod}(S_{g}),\operatorname{Mod}(S_{g})]=\operatorname{Mod}(S_{g}).$
@@
Does the well-suited criterion hold for $\mathcal{H}_{g}$? More precisely, is there a graph $N_{f}$ associated with each $f\in\mathcal{H}_{g}$ so that the connectedness of $N_{f}$ implies that $\mathcal{H}_{g},\mathcal{H}_{g}]$ is contained in the normal closure of $f$ in $\mathcal{H}_{g}$?